OpenStudy (anonymous):

Which of the following has a slope of –2 and a y-intercept of 4? A. y = –2x – 4 B. y = 2x – 4 C. y = 2x + 4 D. y = –2x + 4

OpenStudy (anonymous):

i say that its C

OpenStudy (anonymous):

lol your wrong..... slope is -2 and y-intercept is 4 : y=-2+4

OpenStudy (anonymous):

I agree...in y = mx + b form, the m is the slope and the b is the y intercept....krissy is correct....y = -2x + 4

OpenStudy (solomonzelman):

y=mx+b m- slope b- y-intercept if slope is -2 and y-intercept is -4 then y=(-2)x+(-4), or y=--2x-4

OpenStudy (solomonzelman):

Oh, positive 4, my bad.

OpenStudy (solomonzelman):